14.5 mL of 0.30M NaOH is titrated with 20.2 mL of 0.25M HCl.

What is the final pH of the solution?
Answer #1

NaOH + HCl NaCl + H2O

Millimole of NaOH = molarity × volume (ml) = 0.30 × 14.5 = 4.35

Millimole of HCl = 0.25 × 20.2 = 5.05

Millimole of HCl left = 5.05 - 4.35 = 0.7

Total volume of solution = 14.5ml + 20.2ml = 34.7ml

Molarity of HCl = 0.7 mmol/34.7ml = 0.0201729M

pH = - log[H+] = - log (0.0201729) = 1.695 (answer)